# Speech-Copilot: Leveraging Large Language Models for Speech Processing via Task Decomposition, Modularization, and Program Generation

### v1

@@ Abstract
In this work, we introduce Speech-Copilot, a modular framework for instruction-oriented speech-processing tasks that minimizes human effort in toolset construction. Unlike end-to-end methods using large audio-language models, Speech-Copilot builds speech processing-specific toolsets by analyzing pre-collected task instructions and breaking tasks into manageable sub-tasks. It features a flexible agent based on large language models that performs tasks through program generation. Our approach achieves state-of-the-art performance on the Dynamic-SUPERB benchmark, demonstrating its effectiveness across diverse speech-processing tasks. Key contributions include: 1) developing an innovative framework for speech processing-specific toolset construction, 2) establishing a high-performing agent based on large language models, and 3) offering a new perspective on addressing challenging instruction-oriented speech-processing tasks. Without additional training processes required by end-to-end approaches, our method provides a flexible and extendable solution for a wide range of speech-processing applications.
@@ 1 Introduction
An LLM-based agent capable of utilizing these modules via programming has been developed. Our results show that the developed agent can solve various tasks by appropriately combining the basic modules, achieving state-of-the-art performance on Dynamic-SUPERB [20] compared with baselines including large audio-language models and cascaded systems. This validates the efficacy of Speech-Copilot. We also find that Speech-Copilot has strong multi-task ability that can deal with several tasks in a single user query without sacrificing the performances. A demo page is available111Demo at: https://sites.google.com/view/slt2024-demo-page. After review, we will release all the related code for Speech-Copilot for the community to use, hoping to make it more convenient for everyone to perform speech-processing tasks. Overall, our contributions are:
@@ 2.2 Toolkit Applications in Speech Processing
In contrast, Speech-Copilot effectively addresses these issues through a well-crafted toolset construction and program generation approach, where a toolset with fundamental speech modules is first constructed and the agent can then use these modules as basic blocks and dynamically combine them to solve various tasks based on the user query. This ensures the generalizability and versatility. By solving tasks through programming, Speech-Copilot allows users to modify the program according to their preferences, providing a higher degree of flexibility and enabling behavior manipulation to some extent. Furthermore, analyzing the agent’s reasoning steps in the programs enhances interpretability, making it easier for users to understand how solutions are derived.

### v2

@@ Abstract
In this work, we introduce Speech-Copilot, a modular framework for instruction-oriented speech-processing tasks that minimizes human effort in toolset construction. Unlike end-to-end methods using large audio-language models, Speech-Copilot builds speech processing-specific toolsets by analyzing pre-collected task instructions and breaking tasks into manageable sub-tasks. It features a flexible agent that performs tasks through program generation based on large language models. Our approach achieves state-of-the-art performance on the Dynamic-SUPERB benchmark, demonstrating its effectiveness across diverse speech-processing tasks. Key contributions include: 1) developing an innovative framework for speech processing-specific toolset construction, 2) establishing a high-performing agent based on large language models, and 3) offering a new perspective on addressing challenging instruction-oriented speech-processing tasks. Without additional training required by end-to-end approaches, our method provides a flexible and extendable solution for a wide range of speech-processing applications.
@@ 1 Introduction
An LLM-based agent capable of utilizing these modules via programming is developed. Our results show that this agent can solve various tasks by appropriately combining the basic modules, achieving state-of-the-art performance on Dynamic-SUPERB [20] compared with baselines including large audio-language models and cascaded systems. This validates the efficacy of Speech-Copilot. We also find that Speech-Copilot has strong multi-task ability that can deal with several tasks in a single user query without sacrificing the performances. The demo page and code for Speech-Copilot are available111https://github.com/kuan2jiu99/speech-copilot. Overall, our contributions are:
@@ 2.2 Toolkit Applications in Speech Processing
In contrast, Speech-Copilot effectively addresses these issues through a well-crafted toolset construction and program generation approach, where a toolset with fundamental speech modules is first constructed and the agent can then use these modules as basic blocks and dynamically combine them to solve various tasks based on the user query. This ensures the generalizability and versatility. By solving tasks through programming, Speech-Copilot allows users to modify the program according to their preferences, providing a higher degree of flexibility and enabling behavior manipulation to some extent. Furthermore, analyzing the agent’s reasoning steps in the programs enhances interpretability, making it easier for users to understand how the solutions are derived.
@@ 6 Conclusion and Future work
7 Acknowledgement
This work was inspired by a discussion between Prof. Thomas Hain and Prof. Hung-yi Lee during a taxi ride at ASRU 2023. In addition, we thank the National Center for High-performance Computing (NCHC) of the National Applied Research Laboratories (NARLabs) in Taiwan for providing computational and storage resources.